# Winding quotients for virtual period maps of rank 1

### v4

@@ 6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
Yosuke Ohyama has informed author that the function $\mathcal{W}(w,q)$ can be expressed by a bilateral basic hypergeometric series ([GR] chap.5.) as follows;
$_{2}\psi_{2}\begin{bmatrix}\vskip-8.5359ptw,w&\\ \vskip-5.69046pt&\!\!;q,q\\ wq,wq\!\!&\end{bmatrix}=\sum_{n=-\infty}^{\infty}\!\frac{(w,w;q)_{n}}{(wq,wq;q)_{n}}q^{n}=\sum_{n=-\infty}^{\infty}\!\frac{(1-w)^{2}}{(1-wq^{n})^{2}}q^{n}=\frac{(1-w)^{2}}{w}\mathcal{W}(w,q).$
@@ 7. Solution to the new inversion problem
The descriptions of the morphism: $\mathbb{C}_{w}^{\times}\to E_{\tau}\subset\mathbb{P}^{2}$ is achieved as follows. This gives an answer to the new inversion problem posed at the end of section 5.
@@
The pair of meromorphic functions in $w\in\mathbb{C}_{w}^{\times}$
$\begin{array}[]{ccl}(\ 2\pi\sqrt{-1})^{2}\mathcal{W}(w,\exp(2\pi\sqrt{-1}\tau)-\frac{\pi^{2}}{3}E_{2}(\tau)\ :\ (2\pi\sqrt{-1})^{3}\mathcal{W}^{\prime}(w,\exp(2\pi\sqrt{-1}\tau)\ :\ 1\ )\end{array}$
gives a morphism $\mathbb{C}_{w}^{\times}\to E_{\tau}\subset\mathbb{P}^{2}$, making the diagram ()(*) in section 5 commutative.
@@
Acknowledgements. The author express his gratitude to Hiroki Aoki who informed him the Borcherds work (see the proof of Assertion B) and helped him in solving the new inversion problem, and to Akishi Ikeda who explained him the roles of the pull-back of the function $\mathcal{W}$ to the winding covering space in mathematical physics and informed him its related literatures (see Remark 1.1). He also express his gratitude to Masahiko Yoshinaga for discussions during the preparation of the present note.

### v5

@@ 5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
Yosuke Ohyama has informed author the followings
1. The function $\mathcal{U}(w,q)$ gives the double product part of Jacobi triple product:
$\sum_{m=-\infty}^{m=\infty}q^{\frac{m(m-1)}{2}}(-w)^{m}=\prod_{m=1}^{\infty}(1-q^{m})\prod_{m=0}^{\infty}(1-wq^{m})\prod_{m=1}^{\infty}(1-w^{-1}q^{m})$
2. The function $\mathcal{V}_{+}(w,q)$ is expressed by a Lambert series [Kn] p.452 (f) on the disc $\mathbb{D}_{w}$.
$\qquad\qquad\sum_{n=0}^{\infty}\frac{w^{n}q^{n}}{1-q^{n}}\ =\sum_{n=0}^{\infty}\frac{wq^{n}}{1-wq^{n}}\Big|_{\mathbb{D}_{w}\times\mathbb{D}_{q}}\ =\ \mathcal{V}(w,q)\Big|_{\mathbb{D}_{w}\times\mathbb{D}_{q}}\vskip-5.69046pt$
3. The function $\mathcal{W}(w,q)$ is expressed by bilateral basic hypergeometric series ([GR] (chap.5.).
$_{2}\psi_{2}\begin{bmatrix}w,w&\\ \vskip-5.69046pt&\!\!;q,q\\ wq,wq\!\!&\end{bmatrix}=\!\sum_{n=-\infty}^{\infty}\!\frac{(w,w;q)_{n}}{(wq,wq;q)_{n}}q^{n}=\!\sum_{n=-\infty}^{\infty}\!\frac{(1-w)^{2}}{(1-wq^{n})^{2}}q^{n}=\frac{(1-w)^{2}}{w}\mathcal{W}(w,q).$
@@ 7. Solution to the new inversion problem
Using the $q$-multiplicatively periodic function $\mathcal{W}(w,q)$ constructed in the previous section, the descriptions of the morphism: $\mathbb{C}_{w}^{\times}\to E_{\tau}\subset\mathbb{P}^{2}$ is achieved as follows. This gives an answer to the new inversion problem posed at the end of section 5.
@@
For each fixed $q=\exp(2\pi\sqrt{-1}\tau)$, the pair of functions:
$(\ (2\pi\sqrt{-1})^{2}\mathcal{W}(w,q)\!-\!\frac{\pi^{2}}{3}E_{2}(\tau)\ :\ (2\pi\sqrt{-1})^{3}w\frac{\partial}{\partial w}\mathcal{W}(w,q)\ :\ 1\ )$
gives a morphism: $\mathbb{C}_{w}^{\times}\to E_{\tau}\subset\mathbb{P}^{2}$ making the diagram ()(*) in section 5 commutative.
@@
8. Appendix
We describe the relationship between Euler’s symbol $(w;q)_{\infty}$ and the Weierstrass elliptic functions. We use the results and the equations (1) - (10) in section 6.
Recall that, starting from Euler’s symbol $(w;q)_{\infty}=\prod_{m=0}^{\infty}(1-wq^{m})$ (1), we have introduced the (either holomorphic or meromorphic) functions on $(w,q)\in\mathbb{C}_{w}^{\times}\times\mathbb{C}_{q}^{\times}$:
$\begin{array}[]{cccccl}\!\!\!\!\!\!(2)\qquad&\mathcal{U}(w,q)&\!=\!&\frac{(w;q)_{\infty}\cdot(w^{-1},q)_{\infty}}{(1-w)}&\!=\!&\prod_{m=0}^{\infty}(1\!-\!wq^{m})\prod_{m=1}^{\infty}(1\!-\!w^{-1}q^{m})\quad\\ \!\!\!\!\!\!(3)\qquad&\mathcal{V}(w,q)&\!=\!&-w\frac{\partial}{\partial w}\log(\mathcal{U}(w,q))&\!=\!&\sum_{m\in\mathbb{Z}_{\geq 0}}\frac{wq^{m}}{1-wq^{m}}+\sum_{m\in\mathbb{Z}_{<0}}\frac{1}{1-wq^{m}}\\ \!\!\!\!\!\!(7)\qquad&\mathcal{W}(w,q)&\!=\!&w\frac{\partial}{\partial w}\mathcal{V}(w,q)&\!=\!&\sum_{m\in\mathbb{Z}}\frac{wq^{m}}{(1-wq^{m})^{2}}\end{array}$
and have shown the relationship
$(9)\qquad\qquad\qquad\qquad\qquad\wp(u,\tau)\ =\ (2\pi\sqrt{-1})^{2}\ \mathcal{W}(w,q)\ -\ \frac{\pi^{2}}{3}E_{2}(\tau)\quad\quad\quad\quad\quad\qquad\qquad\qquad\qquad\quad$
where the coordinates $(w,q)$ for the function $\mathcal{U}(w,q),\mathcal{V}(w,q)$ and $\mathcal{W}(w,q)$ and the coordinates $(u,\tau)$ for the $\wp$-functions are related by the winding quotient morphism
$(10)\qquad\qquad\qquad\qquad\quad\ w=\exp(2\pi\sqrt{-1}u),\qquad q=\exp(2\pi\sqrt{-1}\tau),\qquad\qquad\qquad\quad$
It is straight forward now to show the following.
Assertion C. Weierstrass $\zeta$-function and $\sigma$-function are described by the functions $\mathcal{V}(w,q)$ and $\mathcal{U}(w,q)$ as follows, where we use the coordinate change (10).
$(11)\qquad\qquad\quad\!\zeta(u,\tau)\ =\ -2\pi\sqrt{-1}\ \mathcal{V}(w,q)\ +\ \frac{\pi^{2}}{3}E_{2}(\tau)u\ -\ \pi\sqrt{-1}\qquad\qquad\quad\quad\quad\ \ \$
$(12)\qquad\quad\quad\sigma(u,\tau)=-\frac{1}{2\pi\sqrt{-1}}\frac{\mathcal{U}(w,q)}{\eta(\tau)^{2}}\exp\Big(\frac{\pi^{2}}{6}E_{2}(\tau)u^{2}-\pi\sqrt{-1}u+\frac{\pi\sqrt{-1}}{6}\tau\Big).\quad\quad\quad$
where $\eta(\tau)$ is Dedekind eta-function.
Proof.
Recall that Weierstrass $\zeta$-function and $\sigma$-function are defined by the equations $\wp(u,\tau)=-\frac{\partial}{\partial u}\zeta(u,\tau)$ and $\frac{\partial}{\partial u}\log\sigma(u,\tau)=\zeta(u,\tau)$ together with some initial conditions. Note also that we have $\frac{\partial}{\partial u}=2\pi\sqrt{-1}w\frac{\partial}{\partial w}$ due to (10).
Case for $\zeta(u,\tau)$.
In view of (7) and (9), we have $\frac{\partial}{\partial u}\zeta(u,\tau)+(2\pi\sqrt{-1})^{2}w\frac{\partial}{\partial w}\mathcal{V}(w,q)-\frac{\pi^{2}}{3}E_{2}(\tau)=0$. We integrate it by $du\!=\!\frac{1}{2\pi\sqrt{-1}}\frac{dw}{w}$, where we set the integral constant to be $F(\tau)$. That is,
$(*)\qquad\qquad\qquad\qquad\zeta(u,\tau)\ +\ 2\pi\sqrt{-1}\mathcal{V}(w,q)\ -\ \frac{\pi^{2}}{3}E_{2}(\tau)u\ =\ F(\tau).\qquad\qquad\qquad\vskip-2.84544pt$
Recall that the zeta function is an odd function w.r.t. the variable $u$. On the other hand, the involution $u\leftarrow\!\!\!\rightarrow-u$ induces the involution $w\leftarrow\!\!\!\rightarrow w^{-1}$. We apply the involutions to ()(*). Recalling the involution relation (5) on $\mathcal{V}(w,q)$, we obtain
$\quad-\zeta(u,\tau)\ -\ 2\pi\sqrt{-1}(\mathcal{V}(w,q)\ +\ 1)\ +\ \frac{\pi^{2}}{3}E_{2}(\tau)u=F(\tau)\vskip-2.84544pt$
Suming this with ()(*), we get $2F(\tau)=-2\pi\sqrt{-1}$, and the formula (11).
Case for $\sigma(w,\tau)$.
In view of (3) and (11), we have $\frac{\partial}{\partial u}\log\sigma(u,\tau)=2\pi\sqrt{-1}\ w\frac{\partial}{\partial w}\log\mathcal{U}(w,q)+\frac{\pi^{2}}{3}E_{2}(\tau)u-\pi\sqrt{-1}$. We integrate it again by $du=\frac{1}{2\pi\sqrt{-1}}\frac{dw}{w}$, where we set the multiplicative integral constant to be $G(\tau)$. That is,
$\qquad\qquad\qquad\qquad\ \ \sigma(u,\tau)=\mathcal{U}(w,q)\exp\big(\frac{\pi}{6}E_{2}(\tau)u^{2}-\pi\sqrt{-1}u\big)G(\tau)\qquad\qquad\qquad\ \ \vskip-2.84544pt$
Recall that $\underset{u\to 0}{\lim}\frac{\sigma(u,q)}{u}=1$, where $u\to 0$” is the same as $w\to 1$”. Thus, we obtain
$1=\underset{w\to 1}{\lim}\frac{2\pi\sqrt{-1}\mathcal{U}(w,q)}{\log(w)}\cdot G(\tau)\vskip-2.84544pt$
where we take the principal value of $\log(w)$. Recall the Tayler expansion $\log(w)=(w-1)-\frac{1}{2}(w-1)^{2}+\cdots$ at $w=1$. Using (2), we obtain
$2\pi\sqrt{-1}\frac{\mathcal{U}(w,1)}{w-1}\Big|_{w=1}=-2\pi\sqrt{-1}\prod_{m>0}(1-q^{m})^{2}=-2\pi\sqrt{-1}\exp(-\pi\sqrt{-1}\tau/6)\eta(\tau)^{2},\vskip-8.5359pt$
so that $G(\tau)=-\frac{\exp(\frac{\pi\sqrt{-1}}{6}\tau)}{2\pi\sqrt{-1}\eta(\tau)^{2}}$. So, we get the formula (12). ∎
Note. That $\sigma(w,\tau)$ is an odd function in $u$ follows immediately from (12) and (5). However it does not determine the constant term $\frac{\pi\sqrt{-1}}{6}\tau$ in the exponential part of (12).
@@ 8. Appendix
Acknowledgements. The author express his gratitude to Hiroki Aoki who informed him the Borcherds work (see the proof of Assertion B) and helped him in solving the new inversion problem, to Akishi Ikeda who explained him the roles of the pull-back of the function $\mathcal{W}$ to the winding covering space in mathematical physics and informed him its literatures (see Remark 1.1), and to Yosuke Ohyama for guiding to bilateral hypergeometric functions and related interesting discussions on references and histories (see Remark 6.3). He also express his gratitude to Masahiko Yoshinaga for discussions during the preparation of the present note.